OpenStudy (anonymous):

Use implicit differentiation to find the slope of the tangent line to the curve -3x^2 - 1xy -2y^3 = -54 at the point (-1, 3) I am really frustrated with this

OpenStudy (amistre64):

show mw what youve tried doing

OpenStudy (anonymous):

The derivative I got was y' = -3x^2 / -x-y-6y^2

OpenStudy (anonymous):

The derivative of -3x^2 is -6x

OpenStudy (anonymous):

I used product rule for -1xy

OpenStudy (amistre64):

and -xy ?

OpenStudy (anonymous):

and i got -y + -xy

OpenStudy (amistre64):

nope

OpenStudy (anonymous):

ah

OpenStudy (anonymous):

so what would it be?

OpenStudy (amistre64):

-y-xy'

OpenStudy (anonymous):

ok so we are at -6x - y -xy'?

OpenStudy (amistre64):

product rule is: -x'y - xy' ; since x' = 1 , we can simplify to -y-xy'

OpenStudy (amistre64):

so far yes

OpenStudy (anonymous):

Ok and the derivative of -2y^3 is -6y^2y' , and the 54 would be trivial because when derivatives are taken it would be canceled?

OpenStudy (amistre64):

yes and the 54 derives to 0 like all constants do

OpenStudy (anonymous):

ok so now I would want to get the y' by itself

OpenStudy (amistre64):

you can either plug in for x and y as is and solve for y' afterwards; or you can get y' by itself and plug in for x and y ... either way

OpenStudy (anonymous):

Oh i see

OpenStudy (anonymous):

Let me work this out and give you what I got?

OpenStudy (anonymous):

sorry if i sound dumb

OpenStudy (anonymous):

Ok are u still there?

OpenStudy (anonymous):

I just want to make sure I have it right

OpenStudy (amistre64):

im here

OpenStudy (anonymous):

ok so i am at -6(-1)-3 + 1y' -6(3)^2y'

OpenStudy (anonymous):

using the plug in method that you said

OpenStudy (amistre64):

looks good so far except you left off the =0 :)

OpenStudy (anonymous):

damn i am having trouble with this plugging in with the y' man i am so dumb...

OpenStudy (amistre64):

6-3+y'-54y' = 0 3+y'(1-54) = 0 y' = 3/53

OpenStudy (anonymous):

ok great but the 54?

OpenStudy (anonymous):

you subtracted oh ok i see

OpenStudy (amistre64):

:) you have to be able to retain the algebra stuff while learning the calculus stuff

OpenStudy (anonymous):

yeah true thanks a lot man :)

OpenStudy (amistre64):

good luck ;)

OpenStudy (anonymous):

u saved me :)

OpenStudy (anonymous):

and thanks